Provide the values of β that are the solutions to the equation:

2cos(3β + π/3) = -√2



1
Expert's answer
2021-05-19T14:19:42-0400

Provide the values of β that are the solutions to the equation:

"2\cos\left(3\beta+\frac{\pi}{3}\right)=-\sqrt2".

"\textbf{Solution:}"

"2\cos\left(3\beta+\frac{\pi}{3}\right)=-\sqrt2".


"\left[ 
\begin{gathered} 
3\beta+\frac{\pi}{3}=\cos^{-1}\left(-\frac{\sqrt2}{2}\right)+2\pi k, \, k\in\mathbb {Z}, 
\\
3\beta+\frac{\pi}{3}=-\cos^{-1}\left(-\frac{\sqrt2}{2}\right)+2\pi k, \, k\in\mathbb {Z}. 
\end{gathered}
\right."

"\left[ 
\begin{gathered} 
3\beta+\frac{\pi}{3}=\frac{3\pi}{4}+2\pi k, \, k\in\mathbb {Z},
\\
3\beta+\frac{\pi}{3}=-\frac{3\pi}{4}+2\pi k, \, k\in\mathbb {Z}. 
\end{gathered}
\right."

"\left[ 
\begin{gathered} 
3\beta=\frac{5\pi}{12}+2\pi k, \, k\in\mathbb {Z}, 
\\
3\beta=-\frac{13\pi}{12}+2\pi k, \, k\in\mathbb {Z}. 
\end{gathered}
\right."

"\left[ 
\begin{gathered} 
\beta=\frac{5\pi}{36}+\frac{2\pi k}{3}, \, k\in\mathbb {Z}, 
\\
\beta=-\frac{13\pi}{36}+\frac{2\pi k}{3}, \, k\in\mathbb {Z}. 
\end{gathered}
\right."

"\textbf{Answer:}"

"\boxed{\beta\in \{\frac{5\pi}{36}+\frac{2\pi k}{3} \}\cup \{-\frac{13\pi}{36}+\frac{2\pi k}{3}\}, \, k\in\mathbb {Z}}"